### 16 Apr 2026 v1

# Coherent control of optomechanical entanglement and steering via dual parametric amplification

Abstract
We propose a coherent-control scheme for engineering quantum correlations in a cavity optomechanical (COM) system consisting of a driven optical cavity with an embedded nonlinear medium and a membrane, assisted by a coherent feedback loop. The nonlinear medium and the membrane are pumped to implement optical and mechanical parametric amplifications with controllable modulation frequencies and pump amplitudes. Through the combined modulation of the two parametric amplifications and the coherent feedback loop, we engineer the effective cavity decay rate and the distribution of quantum fluctuations, thereby strengthening quantum correlations and improving their robustness against thermal noise. Our scheme provides an efficient route to realizing highly tunable, strong, thermally robust quantum correlations in COM systems, which is promising for the protection of fragile quantum resources.
preprint: APS/123-QED
I Introduction:
Quantum entanglement [1, 2, 3, 4, 5], as one of the most fundamental features of quantum mechanics, is a key source for quantum metrology [6, 7], quantum computation [8, 9], and quantum communication[10, 11]. It has been demonstrated in atoms [12], ions [13], and cavity optomechanical (COM) systems [14, 15, 16, 17], where the interaction between an optical cavity and mechanical motion is mediated by radiation pressure [18, 19]. COM systems have enabled a variety of important phenomena and applications, such as mechanical squeezing [20, 21], optomechanically induced transparency [22, 23], and mechanical resonator cooling [24, 25]. Beyond entanglement, Einstein-Podolsky-Rosen (EPR) steering is a form of quantum nonlocality intermediate between entanglement and Bell nonlocality [26, 27]. Several schemes to generate steady-state one-way steering in COM systems have also been explored [28, 29]. To protect and enhance the fragile quantum correlations in COM platforms, substantial efforts have been devoted to developing various approaches. Among these, parametric driving enabled by optical and mechanical parametric amplification (OPA and MPA) provides a versatile and effective means to enhance nonlinearity and access rich classical and quantum dynamics [30, 31, 32, 33, 34, 35].
Feedback control is also a viable strategy for effectively enhancing quantum correlations in COM systems by extracting and re-injecting quantum information [36, 37, 38, 39, 40]. Broadly, feedback mechanisms can be divided into two categories: measurement-based feedback [41, 42] and coherent feedback [43, 44]. The former is typically complicated, as it relies on measurement and subsequent classical processing, which can introduce additional noise and backaction. In contrast, coherent feedback avoids explicit measurement and does not introduce measurement-induced noise. It is therefore attracted interest for applications such as improving resonator cooling [45], protecting qubit coherence [46], and enhancing entanglement and steering [17, 16].
Inspired by pervious studies, we propose a scheme to enhance optomechanical entanglement and EPR steering in a COM system consisting of a nonlinear medium and a membrane embedded in a Fabry–Pérot (FP) cavity with a feedback loop. The loop is formed by a highly reflect mirror (HRM) and a controllable beam splitter (CBS), which feeds the cavity output field back into the cavity. The nonlinear medium and the membrane are driven by two laser fields to implement OPA and MPA, while an additional external laser field pumps the FP cavity through the CBS. Our proposal combines two effective ways and are expected to amplify quantum correlations. Within experimentally feasible parameter regimes, we first investigate the limit-cycle cavity dynamics and entanglement as functions of the CBS reflectivity. As the reflectivity increases, the effective cooperativity is enhanced and the system approaches the dynamical-instability threshold, leading to a larger cavity oscillation amplitude. Meanwhile, the maximum entanglement exhibits a non-monotonic dependence, increasing at first and then decreasing. This behavior indicates that a moderate reflectivity can be chosen to enable constructive interference of quantum-noise pathways between the output field and the input field, thereby amplifying the entanglement. We then demonstrate the joint modulation induced by parametric amplification and the coherent feedback loop. The maximum entanglement is evaluated as a function of the OPA-MPA modulation-frequency ratio (amplitude ratio) and reflectivity. Using the same set of parameters, we further evaluate the maximum steering and find coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables the transition between one-way and two-way steering in the COM system. By varying the frequency ratio, amplitude ratio, and reflectivity, one can enhance the steering and control its strength as well as the transition between one-way and two-way steering. Finally, we show the quantum correlations become more robust against thermal noise. The results provide physical insight into multi-field driven COM systems with coherent feedback. Our scheme offers an effective route to strong, high-quality, and widely tunable entanglement and steering, with potential applications in quantum computation and quantum communication.
II System and dynamics:
Figure 1: Schematic diagram of the proposed COM system with the embedded membrane and the $\chi^{(2)}$ nonlinear medium, coupled to the coherent feedback loop. (a) The membrane and the nonlinear medium are each pumped by two laser fields with amplitude $G_{m}$ and $G_{c}$, frequency $\omega_{m}$ and $\omega_{c}$, respectively. (b)The coherent feedback loop consists of the HRM and the CBS. The external laser field with amplitude $E$ drives the cavity via the CBS. Here $a_{\mathrm{in}}$ denotes the optical input vacuum noise associated with the zero-point fluctuation entering the cavity through the CBS. The blue arrows indicate that the optical output field $a_{\mathrm{out}}$ reflected by the HRM is fed back into the cavity through the CBS.
As shown in Fig. 1, we consider a COM setup in which a nonlinear $\chi^{(2)}$ medium and a membrane are embedded inside a Fabry–Pérot (FP) cavity with a feedback loop. The FP cavity supports a photon mode with frequency $\omega_{a}$. The membrane acting as a mechanical oscillator supports a phonon mode with effective mass $m$ and frequency $\omega_{b}$ [16]. The cavity is pumped by an external laser field with amplitude $E$, transmitted through a controllable beam splitter (CBS) with transmission $t_{b}$ and reflectivity $r_{b}$, satisfying $t_{b}^{2}+r_{b}^{2}=1$ [17]. $E=\sqrt{2\kappa_{a}P_{l}/(\hbar\omega_{l})}$ is the amplitude of the laser, where $\kappa_{a}$ is the decay rate of the optical cavity, $P_{l}$ ($\omega_{l}$) is the laser power (frequency) [47]. A coherent field with amplitude $G_{c}$, frequency $\omega_{c}$, and phase $\theta_{c}$ is applied to the nonlinear medium inside the cavity, giving rise to optical parametric amplification (OPA) [30, 31]. The membrane is simultaneously subjected to a periodic pump of frequency $\omega_{m}$, amplitude $G_{m}$, and phase $\theta_{m}$, which periodically modulates the spring constant and produces mechanical parametric amplification (MPA) [35, 15]. The Hamiltonian of the COM systems in a frame rotating with respect to the driving laser frequency $\omega_{l}$ reads [30]
${H}/\hbar$ $=\Delta_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-ga^{\dagger}a(b^{\dagger}+b)+\mathrm{i}G_{c}\bigl(a^{\dagger 2}e^{-\mathrm{i}\Delta_{c}t+\mathrm{i}\theta_{c}}-a^{2}e^{\mathrm{i}\Delta_{c}t-\mathrm{i}\theta_{c}}\bigr)$
$\quad+\mathrm{i}G_{m}\bigl(b^{\dagger 2}e^{-\mathrm{i}\omega_{m}t+\mathrm{i}\theta_{m}}-b^{2}e^{\mathrm{i}\omega_{m}t-\mathrm{i}\theta_{m}}\bigr)+\mathrm{i}t_{b}E(a^{\dagger}-a)$ (1)
where $\Delta_{a}=\omega_{a}-\omega_{l}$, $\Delta_{c}=\omega_{c}-2\omega_{l}$. $o(o^{\dagger})(o=a,b)$ are the annihilation (creation) operators of the optical cavity mode and the phonon mode, which satisfy the commutation relation $o,o^{\dagger}]=1$. $g$ is the single-photon optomechanical coupling strength.
In the feedback channel, the output field is reflected by the HRM and then fed back into the optical cavity through the beam splitter. By using the standard input-output relation, the output field is given by $a_{\mathrm{out}}=\sqrt{2\kappa_{a}}a-a_{\mathrm{in}}$ [48]. The feedback-modified input field can be described as a superposition of the original input noise $a_{in}$ and the returned output field. These two contributions are coherently mixed by a lossless CBS before re-entering the cavity, so that the input-field operator can be written as: $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}=r_{b}e^{\mathrm{i}\theta}a_{\mathrm{out}}+t_{b}a_{\mathrm{in}}.$ $\theta$ denotes the phase shift acquired by the output field upon reflection from the HRM. By replacing the original input noise operator $a_{\mathrm{in}}$ by $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}$, the dynamical evolution of the COM system are given by a set of quantum Langevin equations(QLEs):
$\dot{a}$ $=-(\mathrm{i}\Delta_{\mathrm{fb}}+\kappa_{\mathrm{fb}})a+\mathrm{i}g\,a(b^{\dagger}+b)+2G_{c}a^{\dagger}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}+t_{b}E+\sqrt{2\kappa_{a}}\,a_{\mathrm{fb}}^{\mathrm{in}}(t),$ (2)
$\dot{b}$ $=-(\mathrm{i}\omega_{b}+\kappa_{b})b+\mathrm{i}g\,a^{\dagger}a+2G_{m}b^{\dagger}e^{-\mathrm{i}(\omega_{m}t-\theta_{m})}+\sqrt{2\kappa_{b}}\,b^{\mathrm{in}}(t).$
where $\Delta_{\mathrm{fb}}=\Delta_{a}-2\kappa_{a}r_{b}\mathrm{sin}\theta$, $\kappa_{\mathrm{fb}}=\kappa_{a}(1-2r_{b}\mathrm{cos}\theta)$ are respectively the detuning of the photon mode and the effective decay rate. ${a_{\mathrm{fb}}^{\textrm{in}}}(t)=t_{b}(1-r_{b}e^{\mathrm{i}\theta})a^{\mathrm{in}}(t)$. $b^{\text{in}}(t)(\kappa_{b})$ represents the input noise (decay rate) of the phonon mode. The input noise is characterized by the following correlation functions:
$\langle a_{\mathrm{fb}}^{\mathrm{in}}(t)a_{\mathrm{fb}}^{\textrm{in}\dagger}(t^{\prime})$ $=$ $t_{b}^{2}(1-r_{b}e^{\mathrm{i}\theta})(1-r_{b}e^{\mathrm{-i}\theta)}[N_{a}(\omega_{a})+1]\delta(t-t^{\prime}),$
$\langle a_{\mathrm{fb}}^{\textrm{in}\dagger}(t)a_{\mathrm{fb}}^{\textrm{in}}(t^{\prime})\rangle$ $=$ $t_{b}^{2}(1-r_{B}e^{\mathrm{i}\theta})(1-r_{b}e^{\mathrm{-i}\theta)}N_{a}(\omega_{a})\delta(t-t^{\prime}),$
$\langle b^{\mathrm{in}}(t)b^{\mathrm{in}\dagger}(t^{\prime})\rangle$ $=$ $[N_{b}(\omega_{b})+1]\delta(t-t^{\prime}),$
$\langle b^{\mathrm{in}\dagger}(t)b^{\mathrm{in}}(t^{\prime})\rangle$ $=$ $[N_{b}(\omega_{b})+1]\delta(t-t^{\prime}),$ (3)
where $N_{0}(\omega_{0})=[\exp(\frac{\hbar\omega_{0}}{k_{\textrm{B}}T})-1]^{-1}$ is the mean thermal excitation number at the environment temperature $T$, and $k_{\textrm{B}}$ is the Boltzmann constant.
We expand each operator as a sum of its steady-state mean value and a small quantum fluctuation, i.e., $o=\langle o\rangle+\delta o$. By substituting this expression into the QLEs and ignoring the second order fluctuation terms, the linearized QLEs of the quantum fluctuations are given as follows:
$\delta\dot{a}$ $=$ $-(\mathrm{i}\Delta_{\mathrm{fb}}^{\prime}+\kappa_{\mathrm{fb}})\delta a+\mathrm{i}G_{\mathrm{eff}}(\delta b^{\dagger}+\delta b)+2G_{c}\delta a^{\dagger}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}+\sqrt{2\kappa_{a}}{a_{\mathrm{fb}}}^{\textrm{in}}(t),$
$\delta\dot{b}$ $=$ $-(\mathrm{i}\omega_{b}+\kappa_{b})\delta b+\mathrm{i}(G_{\mathrm{eff}}\delta a^{\dagger}+G_{\mathrm{eff}}^{\ast}\delta a)+2G_{m}\delta b^{\dagger}e^{-\mathrm{i}(\omega_{m}t-\theta_{m})}+\sqrt{2\kappa_{b}}{b}^{\textrm{in}}(t),$ (4)
where $\Delta_{\mathrm{fb}}^{\prime}=\Delta_{a}-g(\langle b\rangle+\langle b\rangle^{\ast})-2\kappa_{a}r\mathrm{sin}\theta$, $G_{\mathrm{eff}}=g\langle a\rangle=G_{x}+\mathrm{i}G_{y}$ is the effective COM coupling rate. We can rewrite the linearized QLEs in a matrix form by defining the quadrature fluctuation operators, i.e., $X_{o(o^{\mathrm{in}})}(t)=\frac{1}{\sqrt{2}}[o(o^{\mathrm{in}})+o^{\dagger}(o^{\mathrm{in}\dagger})]$ and $Y_{o(o^{\mathrm{in}})}(t)=\frac{1}{\sqrt{2}\mathrm{i}}[o(o^{\mathrm{in}})-o^{\dagger}(o^{\mathrm{in}\dagger})]$. The matrix form is $\dot{u}(t)=A(t)u(t)+n(t)$, where $u^{T}(t)=[\delta X_{a}(t),\delta Y_{a}(t),\delta X_{b}(t),\delta Y_{b}(t)]$ and $n^{T}(t)=[\sqrt{2\kappa_{a}}X_{a}^{\mathrm{in}},\sqrt{2\kappa_{a}}Y_{a}^{\mathrm{in}},\sqrt{2\kappa_{b}}X_{b}^{\mathrm{in}},\sqrt{2\kappa_{b}}Y_{b}^{\mathrm{in}}]$. The drift matrix $A(t)$ is given as follows:
$A(t)=\left(\begin{array}[]{c c c c }-\kappa_{\mathrm{fb}}+\Gamma_{a}&\Delta_{\mathrm{fb}}^{\prime}-\zeta_{a}&-2G_{y}&0\\ -\Delta_{\mathrm{fb}}^{\prime}-\zeta_{a}&-\kappa_{\mathrm{fb}}-\Gamma_{a}&2G_{x}&0\\ 0&0&-\kappa_{b}+\Gamma_{m}&\omega_{b}-\zeta_{m}\\ 2G_{x}&2G_{y}&-\omega_{m}-\zeta_{m}&-\kappa_{b}-\Gamma_{m}\\ \end{array}\right),$ (5)
where $\Gamma_{m}=2G_{m}\mathrm{cos}(\omega_{m}t-\theta_{m})$, $\Gamma_{a}=2G_{a}\mathrm{cos}(\Delta_{a}t-\theta_{a})$, $\zeta_{m}=2G_{m}\mathrm{sin}(\omega_{m}t-\theta_{m})$, $\zeta_{a}=2G_{a}\mathrm{sin}(\Delta_{a}t-\theta_{a})$. We can find the dynamics of the system can be effected by the effects of OPA, MPA, and the coherent feedback loop. Therefore, we show that the dynamics of the system can be regulated by the reflectivity of the CBS in Fig. 2. The dynamics of the quantum fluctuations can be fully characterized by the covariance matrix (CM) $\mathbf{V}$, a $4\times 4$ real symmetric matrix, with its matrix elements defined as $V_{ij}=\frac{1}{2}\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle$,($i,j=1,2,3,4$). The equation of motion of the CM ($V(t)$) is,
$\dot{V}(t)=\mathbf{A}(t)\mathbf{V}(t)+\mathbf{V}(t)\mathbf{A}(t)^{\mathrm{T}}+\mathbf{D},$ (6)
where $D=\mathrm{diag}[\kappa_{a}t_{b}^{2}(1-r_{b})^{2}(2N_{\mathrm{a}}+1),\kappa_{a}t_{b}^{2}(1-r_{b})^{2}(2N_{\mathrm{a}}+1),\kappa_{b}(2N_{b}+1),\kappa_{b}(2N_{b}+1)]$ is the diffusion matrix. Here we choose $\theta=0$ for simplicity.
The CM can be described by block matrix form
$V(t)=\left(\begin{array}[]{c c }V_{a}&V_{ab}\\ V_{ab}^{T}&V_{b}\\ \end{array}\right),$ (7)
where $V_{a}$, $V_{b}$, $V_{ab}$ are $2\times 2$ subblock matrices corresponding to the photon mode, the phonon mode and their correlation, respectively.
In order to verify the bipartite COM entanglement, we use logarithmic negativity $E_{N}$, which is defined as [49]
$E_{N}=\mathrm{max}[0,-\mathrm{ln}{2\eta^{-}}],$ (8)
where $\eta^{-}=\frac{1}{\sqrt{2}}\sqrt{\Sigma-\sqrt{\Sigma^{2}-4\mathrm{det}(V_{ab})}}$, $\Sigma=\mathrm{det(V_{a})}+\mathrm{det}(V_{b})-2\mathrm{det}V_{ab}$.
The steering from the photon mode to the phonon mode, i.e., the photon mode can steer the phonon mode, is defined as,
$\mathcal{G}_{a\rightarrow b}$ $=$ $\max[0,\frac{1}{2}\mathrm{ln}\frac{\mathrm{det}V_{a}}{4\mathrm{det}V_{ab}}].$ (9)
Similarly, the steering from the phonon mode to the photon mode is given by
$\mathcal{G}_{b\rightarrow a}$ $=$ $\max\left[0,\frac{1}{2}\mathrm{ln}\frac{\mathrm{det}V_{b}}{4\mathrm{det}V_{ab}}\right].$ (10)
When studying the squeezing feature of the phonon mode, the squeezing degree of a quadrature component can be characterized as (in units of $\mathrm{dB}$)
$S(O)=-10\log_{10}\!\left[\frac{\langle\delta O(t)^{2}\rangle}{\langle\delta O(t)^{2}\rangle_{\mathrm{vac}}}\right],$ (11)
where $O=X_{b},Y_{b}$ denote the position and momentum quadratures of the phonon mode, respectively. For a vacuum environment, $\langle\delta O(t)^{2}\rangle_{\mathrm{vac}}=1/2$. To further investigate the squeezing effect of the phonon mode, we consider the reduced covariance matrix $V_{b}$ of the phonon mode. Let $\lambda_{\min}$ denote the smallest eigenvalue of $V_{b}$. Then the squeezing degree of the phonon mode is defined as
$S_{b}=-10\log_{10}\!\left(2\lambda_{\min}\right).$ (12)
In addition, the purity of the phonon mode is defined as
$\mu_{b}=\mathrm{Tr}(\rho_{b}^{2}),$ (13)
where $\rho_{b}$ is the reduced density matrix of the phonon mode. For a single mode Gaussian state, the purity can be equivalently written as
$\mu_{b}=\frac{1}{2\bar{n}_{b}+1},$ (14)
where $\bar{n}_{b}$ denotes the thermal occupation number, namely, the average number of excitations in the number state basis in which $\rho_{b}$ is a thermal state.
According to Eq. (2), the classical mean values $\langle o(t)\rangle$ may evolve toward an asymptotic periodic orbit according to amplitude-modulated driving schemes [50, 51], which has been verified by the Fig. 2. Therefore, the long-time dynamics of the system is $\tau$ periodic. Then we have $A(t+\tau)=A(t)$ and $V(t+\tau)=V(t)$. Similarly, the entanglement, steering, squeezing degree, and purity will acquire the same periodicity in time as $V(t)$. In this case, we quantify them by its maximum in one period, i.e.,
$E_{N,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[E_{N}(t)],$
$\mathcal{G}_{a\rightarrow b,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[\mathcal{G}_{a\rightarrow b}(t)],$
$\mathcal{G}_{b\rightarrow a,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[\mathcal{G}_{b\rightarrow a}(t)],$
$S_{b,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[S_{b}(t)],$
$\mu_{b,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[\mu_{b}(t)].$ (15)
III Numerical simulations and discussions:
In this section, we study how to control quantum correlations by the synergistic effect of OPA, MPA, and coherent feedback based on numerical simulations. We numerically study how coherent feedback and key system parameters affect quantum correlations in the COM system. We first analyze the bipartite entanglement, with particular emphasis on the roles of coherent feedback, the OPA-MPA modulation-frequency ratio, and the OPA-MPA amplitude ratio. We then analyze steering, focusing on its periodic dynamics, its dependence on the CBS reflectivity and on the OPA–MPA modulation-frequency and amplitude ratios. At least, we analyze the robustness of entanglement and steering against thermal noise. We choose experimentally feasible parameters within the range of typical values [15, 31]: $\omega_{b}/2\pi=1~\mathrm{MHz}$, $\kappa_{a}/2\pi=0.5~\mathrm{MHz}$, $\kappa_{b}/2\pi=1~\mathrm{Hz}$, $E/2\pi=60~\mathrm{GHz}$, $\lambda_{l}=1550~\mathrm{nm}$, $P_{l}\approx 2.9~\mathrm{mW},$ $m=150~\mathrm{ng}$, $g/2\pi=4~\mathrm{Hz}$ and $T=20~\mathrm{mK}$. In the process of numerical simulation, we keep all eigenvalues of matrix $A(t)$ with negative real parts during the numerical simulations.
Figure 2: Limit-cycle cavity dynamics and entanglement versus CBS reflectivity $r$. (a) Long-time evolution of the cavity amplitude $|\langle a(t)\rangle|/10^{4}$ for $r_{b}=0$, $0.15$, and $0.35$. (b) Corresponding limit-cycle trajectories in the phase space of $\alpha=\langle a\rangle$. (c) Time evolution of the logarithmic negativity $E_{N}(t)$ in the stationary regime for the same three values of $r_{b}$. (d) Maximum entanglement over one period, $E_{N,\max}$ (left axis), and the real part of the dominant phonon eigenvalue $\mathrm{Re}[\lambda_{m}]$ (right axis) as functions of $r_{b}$. All other parameters are in the main text.
Figures 2(a)–2(d) show the limit-cycle dynamics of the cavity field and optomechanical entanglement as a function of the CBS reflectivity $r_{b}$. Figs. 2(a) and 2(b) show that increasing the CBS reflectivity $r_{b}$ enlarges both the oscillation amplitude and the radius of the cavity field limit cycle. This is because the coherent feedback reduces the effective cavity decay rate to $\kappa_{\mathrm{fb}}=\kappa_{a}(1-2r_{b})$, so that more photons are trapped in the cavity and the radiation-pressure drive on the phonon mode is enhanced. In Figs. 2(c) and 2(d), we investigate the impact of the feedback CBS reflectivity $r_{b}$ on $E_{N,\max}$. Interestingly, $E_{N,\max}$ exhibits a nonmonotonic behavior: it first increases and then decreases as $r_{b}$ increases. The physical origin is as follows. We introduce an effective cooperativity on the limit cycle, $C_{\mathrm{LC}}=4g^{2}\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}/(\kappa_{\mathrm{fb}}\,\kappa_{b})$ (See Appendix. A). This trend can be understood as the result of two competing effects induced by the coherent feedback. On the one hand, increasing $r_{b}$ enhances the effective cooperativity, thereby strengthening the quantum correlations between the photon mode and the phonon mode and leading to an initial growth of $E_{N,\max}$. On the other hand, a larger $r_{b}$ simultaneously drives the system closer to the dynamical-instability threshold, as indicated by $\mathrm{Re}[\lambda_{m}]$. Moreover, the photon re-injection increases the photon number and amplifies noise fluctuations, which reduces the photon-phonon correlations and $E_{N,\max}$.
Figure 3: Maximum entanglement $E_{N,\mathrm{max}}$ as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (b). The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $G_{m}/G_{c}=1.5$ in (a) and $\omega_{m}/\Delta_{c}=1.6$ in (b). The other parameters are the same as those in Fig. 2.
Figs. 3(a) and 3(b) demonstrate the combined modulation effects of the parametric amplifications and the coherent feedback loop. As shown in Fig. 3(a), the entanglement $E_{N,\mathrm{max}}$ exhibits a pronounced maximum when the frequency of the MPA satisfies $\omega_{m}\approx 1.7\Delta_{c}\approx 2\omega_{b}$. In this situation, the MPA term $\mathrm{i}G_{m}(b^{\dagger 2}e^{-\mathrm{i}\omega_{m}t}-b^{2}e^{\mathrm{i}\omega_{m}t})$ becomes time independent in the rotating frame of the phonon mode. This yields an effective degenerate squeezing Hamiltonian $H_{\mathrm{MPA}}^{\mathrm{eff}}=\mathrm{i}G_{m}(b^{\dagger 2}-b^{2})$, which maximize quadrature squeezing and suppress the thermal noise. The condition is a requirement for phase sensitive amplification in a harmonic oscillator and has been widely used in periodically modulated COM schemes to enhance quantum correlations. The effective decay rate decreases as the reflectivity increases, which has an impact on the entanglement of system. For a fixed modulation frequency ratio, it’s a feasible way to optimize the entanglement by tuning the reflectivity to an appropriate value. However, the entanglement is reduced when the reflectivity approaches $0.35$. Besides, it is observed that many stripes are also in Fig. 3(a). Fig. 3(b) illustrates the influence of the amplitude ratio and the reflectivity on the entanglement. In the absence of both coherent feedback and MPA, the entanglement is almost minimal. When we keep the amplitude of OPA fixed, the entanglement can reach a maximum of $0.42$ with the increase of amplitude of MPA. Besides, the impact of coherent feedback not only makes the tunable parametric area larger, but also enable the entanglement achieves a higher maximum value of $0.52$.
Figure 4: Gaussian steering as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a) and (b), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (c) and (d). The white line denotes the boundary of zero steering. The parameters are $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $G_{m}=0.05\,\omega_{b}$ in (a) and (b), and $E/2\pi=60~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $\omega_{m}/\Delta_{c}=1.7$ in (c) and (d). The other parameters are the same as those in Fig. 3.
In Fig. 4, we investigate the influence of the frequency ratio, amplitude ratio, and CBS reflectivity on the steering in the COM system. As shown in Fig. 4(a)-(b), only one-way steering can be achieved without the coherent feedback. The introduction of coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables a reflectivity-controlled transition between one-way and two-way steering in the COM system. To better show the transition, we also plot several slices at fixed $\omega_{m}/\Delta_{c}$. From these slices, one can see that $G_{b\to a,\max}$ remains finite over the whole range of $r_{b}$, while $G_{a\to b,\max}$ emerges only when $r_{b}$ exceeds a threshold value and then increases. Besides, the maximum steering appears near the area at $r_{b}=0.2$ and $\omega_{m}/\Delta_{c}=0.9$$1.3$, where the parameter is close the stability threshold. This behavior can be understood as follows. The reflectivity directly modifies the effective decay rate of the photon mode and its corresponding matrix elements in the diffusion matrix, while the matrix elements of the phonon mode remain unchanged. There is another obvious phenomenon in which the steering suddenly drop to zero as the $\omega_{m}/\Delta_{c}$ approaches $1$. As we mentioned before, this is because the steering criterion is no longer satisfied due to a sudden change which enables the $\det V_{a}/\det V_{ab}<1$ around this parameter region. Similar abrupt behavior also appears in the numerical calculation of entanglement. Figs. 4(c)-(d) show the steering as functions of the CBS reflectivity and the amplitude ratio. It should be noted that the reflectivity is extended up to 0.3 because the stability condition is less restrictive. By varying the ratio $G_{m}/G_{c}$, one can also enhance the steering and achieve the transition between one-way and two-way steering. Overall, Fig. 4 reveals the combined modulation effects of the frequency ratio, the amplitude ratio, and coherent feedback on steering in the COM system. The CBS reflectivity determines asymmetric steering resulted from the asymmetry of the system [52]. These results show that coherent feedback not only enlarges the steering region, but also enables control over the steering strength and type through the tuning of system parameters. In particular, the existence of one-way steering in our proposed COM system may have promising applications in one-sided device-independent quantum key distribution [53, 54].
Figure 5: Purity and squeezing of the phonon mode as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$. Panel (a) shows the purity, and panel (b) shows the squeezing. The parameters are $E/2\pi=50~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, and $G_{m}/G_{c}=1.5$ in both panels. The other parameters are the same as those in Fig. 4.
High purity is a prerequisite for many mechanical quantum effects. For example, back-action-evading measurements or squeezing require the oscillator to be prepared in a high-purity quantum state [53, 54]. Therefore, in order to show the local quantum characteristics of the phonon mode, we also investigate how the purity and squeezing change with the reflectivity and the frequency ratio in Fig. 5. From Fig. 5(a), the purity sustains a relatively high level over almost the whole figure and only shows a small area with low value around $\omega_{m}/\Delta_{c}=1.7$. With coherent feedback, it’s useful to adjust the reflectivity so as to turn the original low value area into high value area. Fig. 5(b) shows a nearly opposite phenomenon, where most of the area has relatively low values, but the squeezing is almost the highest at $\omega_{m}/\Delta_{c}=1.7$ without coherent feedback. This can also be understood for the similar reason discussed as before. When $\omega_{m}/\Delta_{c}\approx 1.7$, $\omega_{m}$ becomes close to $2\omega_{b}$, which makes the enhancement of MPA induced squeezing on the phonon mode more pronounced. In the other high squeezing area, we can adjust the reflectivity to broaden this region. Compared with Fig. 5(a) and Fig. 5(b), we can find that, by introducing coherent feedback and properly tuning the parameters, the local quantum characteristics of the phonon mode can maintain relatively high purity while still exhibiting strong squeezing. This indicates that the present scheme provides favorable initial-state conditions for observing quantum behavior of the COM systems and offers a promising platform for the preparation of non-classical states of motion, with potential applications in quantum metrology and precision sensing [53, 54, 55].
Figure 6: Maximum entanglement and steering as functions of the feedback phase $\theta$ and the modulation-amplitude ratio $G_{m}/G_{c}$. Panels (a), (b) show $E_{N,\max}$ and $G_{a\rightarrow b,\max}$ respectively. The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.2$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.15$, and $\omega_{m}/\Delta_{c}=1.3$ in (b). The parameters are the same as those in Fig. 5.
Figure 7: Thermal robustness of entanglement and steering under coherent feedback. Panels (a), (b), and (c) show $E_{N,\max}$, $G_{a\rightarrow b,\max}$, and $G_{b\rightarrow a,\max}$, respectively, as functions of the environment temperature $T$ and the reflectivity $r_{b}$. The parameters are $E/2\pi=50~\mathrm{GHz}$, $\Delta_{c}=1.18\,\omega_{b}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=1.5$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=3.0$, and $\omega_{m}/\Delta_{c}=1.5$ in (b) and (c). The other parameters are the same Fig. 6.
As shown in Fig. 6, both the entanglement and the steering exhibit a clear periodic dependence on the phase shift $\theta$, indicating that it provides an efficient means of controlling quantum correlations. Moreover, increasing the modulation-amplitude ratio $G_{m}/G_{c}$ further enhances both $E_{N,\max}$ and $G_{b\rightarrow a,\max}$. It is also worth noting that the high-value region of steering is narrower than that of entanglement, which indicates that steering is more sensitive to the feedback phase and the amplitude ratio.
To further study the effect of thermal noise, we plot the entanglement and steering as functions of the environment temperature $T$ in the presence of coherent feedback. Fig. 7 shows that, within the stability region, properly tuning the reflectivity can effectively improve the thermal robustness. However, as can be seen from Fig. 7(a), when the system is tuned excessively close to the stability threshold, the enhancement of reflectivity has a destructive effect on the entanglement.
IV Experimental Feasibility And Applications:
In this section, we discuss the experimental feasibility of the proposed COM system. It consists of the nonlinear medium and the membrane embedded in the FP cavity assisted by the coherent feedback loop. The optical parametric amplification (OPA) and mechanical parametric amplification (MPA) employed in our system provide effective means to enhance the nonlinearity and enrich the quantum dynamics, and can be implemented as follows.
The OPA has been well developed in both theoretical studies and experimental implementations [51, 56, 57, 58, 59, 60, 61, 62]. It can be realized by placing a PPKTP crystal inside the cavity and driving it with a classical pump field [63, 64, 65]. And the effective strength can be tuned by the pump power, whereas the corresponding squeezing phase can be adjusted through the relative phase locking among the pump, signal, and local-oscillator fields [63, 64].
The MPA can then be activated by a periodic piezoelectric or capacitive drive that modulates the effective spring constant of a Si3N4 membrane. Taking the experimental control of the MPA amplitude and phase as an example, Mashaal et al. recently investigated parametric modulation in high-stress Si3N4 membrane resonators using both piezoelectric and capacitive schemes. In the piezoelectric case, the MPA amplitude was tuned by increasing the parametric drive voltage $V_{p}$, with the modulation strength characterized by $g_{s}=V_{p}/V_{t}$. In the capacitive case, the applied voltage $V(t)$ modulated the effective elastic constant $k_{p}$, which in turn generated the parametric amplification. Moreover, the MPA phase was set by the relative phase $\phi_{p}$ between the parametric modulation and the phase-locked reference signal [66, 67].
To detect and verify the entanglement and steering, it is inevitable to measure the elements of reduced covariance matrix [68, 69]. The cavity field quadratures can be directly accessed by homodyne or heterodyne detection of the cavity output [70, 71]. The mechanical quadratures cannot be directly measured but can be inferred with the help of an auxiliary probe cavity mode which is weak enough to be negligible. When the probe field, with frequency $\omega_{p,l}$, is red detuned from the resonance frequency $\omega_{p,c}$ of the auxiliary probe cavity by the mechanical frequency $\omega_{b}$, i.e., $\omega_{p,l}\approx\omega_{p,c}-\omega_{b}$, the linearized interaction between the probe cavity mode and the mechanical mode reduces to the beam-splitter form, which maps the mechanical state onto the probe output field. In this way, the CM elements associated with the mechanical mode can be reconstructed from the correlations of the probe output field [69, 72, 73].
Besides, the feedback delay $t_{d}$ is an important factor that should be addressed. As we mentioned in Appendix. B, the instantaneous feedback approximation is valid when $|2\kappa_{a}r_{b}t_{d}|\ll 1$. For a feedback loop of length about $10~\mathrm{cm}$, the propagation delay is only of order $10^{-10}$$10^{-9}~\mathrm{s}$, which is much shorter than the cavity field lifetime $1/\kappa_{a}$. Therefore, the feedback delay can be safely neglected, and the coherent feedback can be treated as instantaneous. Also, if the delay is not negligible, imprinting a phase on light in the feedback loop is a feasible approach to cancel the effect of delay [44].
V Conclusion:
In conclusion, we have investigated the role of coherent feedback in regulating the quantum correlations and the local mechanical quantum properties. Due to introducing coherent feedback, the effective decay rate and the corresponding matrix elements of the photon mode in the diffusion matrix are modulated. The coherent feedback in our system can enhance the intracavity photon and make the entanglement stronger and its parameter region broader with the joint modulation of OPA and MPA. Besides, the coherent feedback enlarges the parameter region supporting one way steering and enables the transition from one-way to two-way steering. Moreover, the system can sustain relatively high purity while exhibiting strong squeezing, thereby improving the quality of the local nonclassical resource state. Besides, we also find that compared to the system only modulated by OPA and MPA, our entanglement and steering is more robust against the thermal noises. Our work may provide a more controllable quantum resource, which have potential applications in continuous-variable quantum communication and quantum key distribution.
VI Acknowledgments:
Mei Zhang thanks Junzhong Yang for helpful discussions. Jinhao Jia thanks Lijun He for helpful discussions. Yingru Li thanks Xinchun Li and Juan Yao for helpful discussions. This work was supported by the National Natural Science Foundation of China under Grant No. 11475021 and the National Key Basic Research Program of China under Grant No. 2013CB922000.
Appendix A The limit-cycle dynamics
In this appendix we provide a quantitative explanation of the limit-cycle behavior and the nonmonotonic dependence of the optomechanical entanglement on the CBS reflectivity $r_{b}$ shown in Fig. 2.
Taking the expectation values of the QLEs and neglecting the zero-mean noise terms, we obtain the mean-field equation as follows,
$\dot{\alpha}=-\bigl(\kappa_{\mathrm{fb}}+i\Delta_{\mathrm{fb}}\bigr)\alpha+ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+tE,$ (16)
where we denote $\alpha\equiv\langle a\rangle$ and $\beta\equiv\langle b\rangle$. Then we can obtain the evolution of the photon number $n_{c}(t)=|\alpha(t)|^{2}$ as
$\frac{d}{dt}|\alpha|^{2}=\alpha^{\ast}\dot{\alpha}+\alpha\dot{\alpha}^{\ast}=-2\kappa_{\mathrm{fb}}(r)|\alpha|^{2}+2\,\mathrm{Re}\!\left[\alpha^{\ast}\Bigl(ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+t_{b}E\Bigr)\right].$ (17)
In the long-time regime the system settles into a time-periodic limit cycle. Averaging Eq. (17) over one period $T$ of the limit cycle, we obtain the power-balance relation
$2\kappa_{\mathrm{fb}}(r)\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}=P_{\mathrm{om}}(r)+P_{\mathrm{OPA}}(r)+P_{\mathrm{drv}},$ (18)
where the right-hand side corresponds to the cycle-averaged power injected into the cavity by the optomechanical interaction, by the OPA and by the external drive, respectively.
For the parameter set used in Fig. 2 we have $g,G_{c}\ll\kappa_{\mathrm{fb}}$, so that the drive term $P_{\mathrm{drv}}$ dominates, while $P_{\mathrm{om}}$ and $P_{\mathrm{OPA}}$ provide only small corrections and depend only weakly on $r$. In this experimentally relevant limit Eq. (18) yields the approximate scaling
$\bigl\langle|\alpha|\bigr\rangle_{T}\propto\frac{1}{\kappa_{\mathrm{fb}}(r)}=\frac{1}{\kappa(1-2r\cos\theta)},\qquad$ (19)
which explains why both the oscillation amplitude in Fig. 2(a) and the radius of the limit-cycle trajectories in Fig. 2(b) increase monotonically with the reflectivity $r$.
Appendix B Validity of the instantaneous-feedback approximation
To justify the instantaneous-feedback approximation used in the main text, we explicitly restore the finite propagation delay in the coherent-feedback loop. To avoid confusion with the Floquet period, we denote the feedback delay time by $t_{d}$ and the loop phase by $\phi_{f}$. The feedback-modified input field is then written as
$\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}(t)=r_{b}e^{\mathrm{i}\phi_{f}}a_{\mathrm{out}}(t-t_{d})+t_{b}a_{\mathrm{in}}(t).$ (20)
Using the standard input-output relation
$a_{\mathrm{out}}(t-t_{d})=\sqrt{2\kappa_{a}}\,a(t-t_{d})-a_{\mathrm{in}}(t-t_{d}),$ (21)
the linearized Langevin equation for the cavity fluctuation becomes
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (22)
$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\,\delta a(t-t_{d})+\mathrm{noise},$
where
$\Delta_{a}^{\prime}=\Delta_{a}-g\bigl(\langle b\rangle+\langle b\rangle^{\ast}\bigr).$ (23)
The delay enters only through the feedback term proportional to $\delta a(t-t_{d})$.
When the loop delay is sufficiently short, one may expand the delayed cavity fluctuation as
$\delta a(t-t_{d})=\delta a(t)-t_{d}\,\delta\dot{a}(t)+\mathcal{O}(t_{d}^{2}).$ (24)
Substituting Eq. (24) into Eq. (22) yields
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (25)
$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\left[\delta a(t)-t_{d}\,\delta\dot{a}(t)\right]+\mathrm{noise}.$
Rearranging the terms gives
$\left[1+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}t_{d}\right]\delta\dot{a}(t)$ $=\left[-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\right]\delta a(t)$ (26)
$\quad+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+\mathrm{noise}.$
Therefore, the first-order expansion is valid provided that the prefactor of $\delta\dot{a}(t)$ remains close to unity, namely,
$\left|2\kappa_{a}r_{b}t_{d}\right|\ll 1.$ (27)
This condition gives the criterion under which the propagation delay can be safely neglected and the coherent feedback can be treated as instantaneous.

### 18 Apr 2026 v2

@@
 Our scheme provides an efficient route to realizing highly tunable, strong, thermally robust quantum correlations in COM systems, which is promising for the protection of fragile quantum resources.
-preprint: APS/123-QED
-I Introduction:
+keywords:
+optomechanics, quantum correlations, coherent feedback, parametric amplification
+articletype: Paper
+1 Introduction:
 Quantum entanglement [1, 2, 3, 4, 5], as one of the most fundamental features of quantum mechanics, is a key source for quantum metrology [6, 7], quantum computation [8, 9], and quantum communication[10, 11].
-It has been demonstrated in atoms [12], ions [13], and cavity optomechanical (COM) systems [14, 15, 16, 17], where the interaction between an optical cavity and mechanical motion is mediated by radiation pressure [18, 19].
+It has been demonstrated in atoms [12], ions [13], and cavity optomechanical (COM) systems [14, 15, 16, 17, 18], where the interaction between an optical cavity and mechanical motion is mediated by radiation pressure [19, 20].
 COM systems have enabled a variety of important phenomena and applications, such as mechanical squeezing [21, 22], optomechanically induced transparency [23, 24], and mechanical resonator cooling [25, 26].
@@
 It is therefore attracted interest for applications such as improving resonator cooling [46], protecting qubit coherence [47], and enhancing entanglement and steering [17, 16].
-Inspired by pervious studies, we propose a scheme to enhance optomechanical entanglement and EPR steering in a COM system consisting of a nonlinear medium and a membrane embedded in a Fabry–Pérot (FP) cavity with a feedback loop.
+Inspired by previous studies, we propose a scheme to control optomechanical entanglement and EPR steering in a COM system consisting of a nonlinear medium and a membrane embedded in a Fabry–Pérot (FP) cavity with a feedback loop.
 The loop is formed by a highly reflect mirror (HRM) and a controllable beam splitter (CBS), which feeds the cavity output field back into the cavity.
@@
 Our scheme offers an effective route to strong, high-quality, and widely tunable entanglement and steering, with potential applications in quantum computation and quantum communication.
-II System and dynamics:
+2 System and dynamics:
 Figure 1: Schematic diagram of the proposed COM system with the embedded membrane and the $\chi^{(2)}$ nonlinear medium, coupled to the coherent feedback loop.
-(a) The membrane and the nonlinear medium are each pumped by two laser fields with amplitude $G_{m}$ and $G_{c}$, frequency $\omega_{m}$ and $\omega_{c}$, respectively.
+(a) The membrane and the nonlinear medium are each pumped by two laser fields with amplitudes $G_{m}$ and $G_{c}$, frequencies $\omega_{m}$ and $\omega_{c}$, respectively.
 (b) The coherent feedback loop consists of the HRM and the CBS.
@@
 The membrane is simultaneously subjected to a periodic pump of frequency $\omega_{m}$, amplitude $G_{m}$, and phase $\theta_{m}$, which periodically modulates the spring constant and produces mechanical parametric amplification (MPA) [36, 15].
-The Hamiltonian of the COM systems in a frame rotating with respect to the driving laser frequency $\omega_{l}$ reads [30]
+The Hamiltonian of the COM system in a frame rotating with respect to the driving laser frequency $\omega_{l}$ reads [31]
 ${H}/\hbar$ $=\Delta_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-ga^{\dagger}a(b^{\dagger}+b)+\mathrm{i}G_{c}\bigl(a^{\dagger 2}e^{-\mathrm{i}\Delta_{c}t+\mathrm{i}\theta_{c}}-a^{2}e^{\mathrm{i}\Delta_{c}t-\mathrm{i}\theta_{c}}\bigr)$
@@
 These two contributions are coherently mixed by a lossless CBS before re-entering the cavity, so that the input-field operator can be written as: $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}=r_{b}e^{\mathrm{i}\theta}a_{\mathrm{out}}+t_{b}a_{\mathrm{in}}.$ $\theta$ denotes the phase shift acquired by the output field upon reflection from the HRM.
-By replacing the original input noise operator $a_{\mathrm{in}}$ by $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}$, the dynamical evolution of the COM system are given by a set of quantum Langevin equations(QLEs):
+By replacing the original input noise operator $a_{\mathrm{in}}$ by $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}$, the dynamical evolution of the COM system is given by a set of quantum Langevin equations (QLEs):
 $\dot{a}$ $=-(\mathrm{i}\Delta_{\mathrm{fb}}+\kappa_{\mathrm{fb}})a+\mathrm{i}g\,a(b^{\dagger}+b)+2G_{c}a^{\dagger}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}+t_{b}E+\sqrt{2\kappa_{a}}\,a_{\mathrm{fb}}^{\mathrm{in}}(t),$ (2)
@@
 where $\Gamma_{m}=2G_{m}\mathrm{cos}(\omega_{m}t-\theta_{m})$, $\Gamma_{a}=2G_{a}\mathrm{cos}(\Delta_{a}t-\theta_{a})$, $\zeta_{m}=2G_{m}\mathrm{sin}(\omega_{m}t-\theta_{m})$, $\zeta_{a}=2G_{a}\mathrm{sin}(\Delta_{a}t-\theta_{a})$.
-We can find the dynamics of the system can be effected by the effects of OPA, MPA, and the coherent feedback loop.
+Equation 5 the dynamics of the system can be affected by the effects of OPA, MPA, and the coherent feedback loop.
 Therefore, we show that the dynamics of the system can be regulated by the reflectivity of the CBS in Fig.
@@
 $\mathcal{G}_{b\rightarrow a}$ $=$ $\max\left[0,\frac{1}{2}\mathrm{ln}\frac{\mathrm{det}V_{b}}{4\mathrm{det}V_{ab}}\right].$ (10)
-When studying the squeezing feature of the phonon mode, the squeezing degree of a quadrature component can be characterized as (in units of $\mathrm{dB}$)
+When studying the squeezing feature of the phonon mode, the squeezing degree of a quadrature component can be characterized as [51] (in units of $\mathrm{dB}$)
 $S(O)=-10\log_{10}\!\left[\frac{\langle\delta O(t)^{2}\rangle}{\langle\delta O(t)^{2}\rangle_{\mathrm{vac}}}\right],$ (11)
@@
 $S_{b}=-10\log_{10}\!\left(2\lambda_{\min}\right).$ (12)
-In addition, the purity of the phonon mode is defined as
+In addition, the purity of the phonon mode is defined as [52]
 $\mu_{b}=\mathrm{Tr}(\rho_{b}^{2}),$ (13)
@@
 $\mu_{b,\mathrm{max}}$ $=$ $\underset{\tau}{\max}[\mu_{b}(t)].$ (15)
-III Numerical simulations and discussions:
-In this section, we study how to control quantum correlations by the synergistic effect of OPA, MPA, and coherent feedback based on numerical simulations.
-We numerically study how coherent feedback and key system parameters affect quantum correlations in the COM system.
-We first analyze the bipartite entanglement, with particular emphasis on the roles of coherent feedback, the OPA-MPA modulation-frequency ratio, and the OPA-MPA amplitude ratio.
-We then analyze steering, focusing on its periodic dynamics, its dependence on the CBS reflectivity and on the OPA–MPA modulation-frequency and amplitude ratios.
-At least, we analyze the robustness of entanglement and steering against thermal noise.
+3 Numerical simulations and discussions
+In this section, we study how to control quantum correlations by the synergistic effect of the OPA, MPA, and coherent feedback based on numerical simulations.
+We numerically study how coherent feedback and the OPA-MPA modulation frequency (amplitude) ratio affect quantum correlations in the COM system.
+Moreover, we study the local mechanical quantum property of the phonon mode.
+At last, we analyze the modulation of the phase shift and the robustness of entanglement and steering against thermal noise.
 We choose experimentally feasible parameters within the range of typical values [15, 32]: $\omega_{b}/2\pi=1~\mathrm{MHz}$, $\kappa_{a}/2\pi=0.5~\mathrm{MHz}$, $\kappa_{b}/2\pi=1~\mathrm{Hz}$, $E/2\pi=60~\mathrm{GHz}$, $\lambda_{l}=1550~\mathrm{nm}$, $P_{l}\approx 2.9~\mathrm{mW},$ $m=150~\mathrm{ng}$, $g/2\pi=4~\mathrm{Hz}$ and $T=20~\mathrm{mK}$.
@@
 Figure 3: Maximum entanglement $E_{N,\mathrm{max}}$ as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (b).
-The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $G_{m}/G_{c}=1.5$ in (a) and $\omega_{m}/\Delta_{c}=1.6$ in (b).
+Here, $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $G_{m}/G_{c}=1.5$ in (a) and $\omega_{m}/\Delta_{c}=1.6$ in (b).
 The other parameters are the same as those in Fig.
@@
 In this situation, the MPA term $\mathrm{i}G_{m}(b^{\dagger 2}e^{-\mathrm{i}\omega_{m}t}-b^{2}e^{\mathrm{i}\omega_{m}t})$ becomes time independent in the rotating frame of the phonon mode.
-This yields an effective degenerate squeezing Hamiltonian $H_{\mathrm{MPA}}^{\mathrm{eff}}=\mathrm{i}G_{m}(b^{\dagger 2}-b^{2})$, which maximize quadrature squeezing and suppress the thermal noise.
+This yields an effective degenerate squeezing Hamiltonian $H_{\mathrm{MPA}}^{\mathrm{eff}}=\mathrm{i}G_{m}(b^{\dagger 2}-b^{2})$, which maximizes quadrature squeezing and suppress the thermal noise [55].
 The condition is a requirement for phase sensitive amplification in a harmonic oscillator and has been widely used in periodically modulated COM schemes to enhance quantum correlations.
@@
 However, the entanglement is reduced when the reflectivity approaches $0.35$.
-Besides, it is observed that many stripes are also in Fig.
-3(a).
 Fig.
@@
 When we keep the amplitude of OPA fixed, the entanglement can reach a maximum of $0.42$ with the increase of amplitude of MPA.
-Besides, the impact of coherent feedback not only makes the tunable parametric area larger, but also enable the entanglement achieves a higher maximum value of $0.52$.
+Besides, the impact of coherent feedback not only enlarges the tunable parametric region, but also enables the entanglement achieves a higher maximum value of $0.52$.
 Figure 4: Gaussian steering as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a) and (b), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (c) and (d).
 The white line denotes the boundary of zero steering.
-The parameters are $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $G_{m}=0.05\,\omega_{b}$ in (a) and (b), and $E/2\pi=60~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $\omega_{m}/\Delta_{c}=1.7$ in (c) and (d).
+Here, $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $G_{m}=0.05\,\omega_{b}$ in (a) and (b), and $E/2\pi=60~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $\omega_{m}/\Delta_{c}=1.7$ in (c) and (d).
 The other parameters are the same as those in Fig.
@@
 From these slices, one can see that $G_{b\to a,\max}$ remains finite over the whole range of $r_{b}$, while $G_{a\to b,\max}$ emerges only when $r_{b}$ exceeds a threshold value and then increases.
-Besides, the maximum steering appears near the area at $r_{b}=0.2$ and $\omega_{m}/\Delta_{c}=0.9$$1.3$, where the parameter is close the stability threshold.
+Besides, the maximum steering appears near the area at $r_{b}=0.2$ and $\omega_{m}/\Delta_{c}=0.9$$1.3$, where the parameter is close to the stability threshold.
 This behavior can be understood as follows.
 The reflectivity directly modifies the effective decay rate of the photon mode and its corresponding matrix elements in the diffusion matrix, while the matrix elements of the phonon mode remain unchanged.
-There is another obvious phenomenon in which the steering suddenly drop to zero as the $\omega_{m}/\Delta_{c}$ approaches $1$.
-As we mentioned before, this is because the steering criterion is no longer satisfied due to a sudden change which enables the $\det V_{a}/\det V_{ab}<1$ around this parameter region.
+There is another obvious phenomenon in which the steering suddenly drops to zero as the $\omega_{m}/\Delta_{c}$ approaches $1$.
+As we mentioned before, this is because the steering criterion is no longer satisfied due to a sudden change that makes the $\det V_{a}/\det V_{ab}<1$ around this parameter region.
 Similar abrupt behavior also appears in the numerical calculation of entanglement.
@@
 4 reveals the combined modulation effects of the frequency ratio, the amplitude ratio, and coherent feedback on steering in the COM system.
-The CBS reflectivity determines asymmetric steering resulted from the asymmetry of the system [52].
+The CBS reflectivity determines asymmetric steering resulting from the asymmetry of the system [56].
 These results show that coherent feedback not only enlarges the steering region, but also enables control over the steering strength and type through the tuning of system parameters.
@@
 Panel (a) shows the purity, and panel (b) shows the squeezing.
-The parameters are $E/2\pi=50~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, and $G_{m}/G_{c}=1.5$ in both panels.
+Here, $E/2\pi=50~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, and $G_{m}/G_{c}=1.5$ in both panels.
 The other parameters are the same as those in Fig.
@@
 From Fig.
-5(a), the purity sustains a relatively high level over almost the whole figure and only shows a small area with low value around $\omega_{m}/\Delta_{c}=1.7$.
-With coherent feedback, it’s useful to adjust the reflectivity so as to turn the original low value area into high value area.
+5(a), the purity remains at a relatively high level over almost the entire parameter region and only shows a small area with low value around $\omega_{m}/\Delta_{c}=1.7$.
+With coherent feedback, the reflectivity can be tuned so as to turn the original low value area into high value area.
 Fig.
 5(b) shows a nearly opposite phenomenon, where most of the area has relatively low values, but the squeezing is almost the highest at $\omega_{m}/\Delta_{c}=1.7$ without coherent feedback.
-This can also be understood for the similar reason discussed as before.
-When $\omega_{m}/\Delta_{c}\approx 1.7$, $\omega_{m}$ becomes close to $2\omega_{b}$, which makes the enhancement of MPA induced squeezing on the phonon mode more pronounced.
+This can also be understood for the similar reason discussed above.
+When $\omega_{m}/\Delta_{c}\approx 1.7$, $\omega_{m}$ becomes close to $2\omega_{b}$, which makes the enhancement of MPA-induced squeezing on the phonon mode more pronounced.
 In the other high squeezing area, we can adjust the reflectivity to broaden this region.
-Compared with Fig.
+By comparing Fig.
 5(a) and Fig.
@@
 This indicates that the present scheme provides favorable initial-state conditions for observing quantum behavior of the COM systems and offers a promising platform for the preparation of non-classical states of motion, with potential applications in quantum metrology and precision sensing [57, 58, 59].
-Figure 6: Maximum entanglement and steering as functions of the feedback phase $\theta$ and the modulation-amplitude ratio $G_{m}/G_{c}$.
+Figure 6: Maximum entanglement and steering as functions of the phase shift $\theta$ and the amplitude ratio $G_{m}/G_{c}$.
 Panels (a), (b) show $E_{N,\max}$ and $G_{a\rightarrow b,\max}$ respectively.
-The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.2$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.15$, and $\omega_{m}/\Delta_{c}=1.3$ in (b).
+Here, $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.2$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.15$, and $\omega_{m}/\Delta_{c}=1.3$ in (b).
 The parameters are the same as those in Fig.
@@
 Panels (a), (b), and (c) show $E_{N,\max}$, $G_{a\rightarrow b,\max}$, and $G_{b\rightarrow a,\max}$, respectively, as functions of the environment temperature $T$ and the reflectivity $r_{b}$.
-The parameters are $E/2\pi=50~\mathrm{GHz}$, $\Delta_{c}=1.18\,\omega_{b}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=1.5$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=3.0$, and $\omega_{m}/\Delta_{c}=1.5$ in (b) and (c).
+Here, $E/2\pi=50~\mathrm{GHz}$, $\Delta_{c}=1.18\,\omega_{b}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=1.5$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=3.0$, and $\omega_{m}/\Delta_{c}=1.5$ in (b) and (c).
 The other parameters are the same Fig.
@@
 7(a), when the system is tuned excessively close to the stability threshold, the enhancement of reflectivity has a destructive effect on the entanglement.
-IV Experimental Feasibility And Applications:
+Compared with the scheme in Ref.
+[15], the quantum correlations are more robust against thermal noise.
+4 Experimental Feasibility And Applications
 In this section, we discuss the experimental feasibility of the proposed COM system.
@@
 It can be realized by placing a PPKTP crystal inside the cavity and driving it with a classical pump field [67, 68, 69].
-And the effective strength can be tuned by the pump power, whereas the corresponding squeezing phase can be adjusted through the relative phase locking among the pump, signal, and local-oscillator fields [63, 64].
+The effective strength can be tuned by the pump power, whereas the corresponding squeezing phase can be adjusted through the relative phase locking among the pump, signal, and local-oscillator fields [67, 68].
 The MPA can then be activated by a periodic piezoelectric or capacitive drive that modulates the effective spring constant of a Si3N4 membrane.
@@
 Moreover, the MPA phase was set by the relative phase $\phi_{p}$ between the parametric modulation and the phase-locked reference signal [70, 71].
-To detect and verify the entanglement and steering, it is inevitable to measure the elements of reduced covariance matrix [68, 69].
+To detect and verify the entanglement and steering, it is necessary to measure the elements of the reduced CM[72, 73].
 The cavity field quadratures can be directly accessed by homodyne or heterodyne detection of the cavity output [74, 75].
@@
 Therefore, the feedback delay can be safely neglected, and the coherent feedback can be treated as instantaneous.
-Also, if the delay is not negligible, imprinting a phase on light in the feedback loop is a feasible approach to cancel the effect of delay [44].
-V Conclusion:
-In conclusion, we have investigated the role of coherent feedback in regulating the quantum correlations and the local mechanical quantum properties.
-Due to introducing coherent feedback, the effective decay rate and the corresponding matrix elements of the photon mode in the diffusion matrix are modulated.
-The coherent feedback in our system can enhance the intracavity photon and make the entanglement stronger and its parameter region broader with the joint modulation of OPA and MPA.
-Besides, the coherent feedback enlarges the parameter region supporting one way steering and enables the transition from one-way to two-way steering.
+Also, if the delay is not negligible, imprinting a phase on light in the feedback loop is an feasible approach to cancel the effect of delay [45].
+5 Conclusion
+In conclusion, we investigated the synergistic effect of the coherent feedback and parametric amplifications in regulating the quantum correlations and the local mechanical quantum properties.
+Our investigation shows that the coherent feedback has two roles: the modulation of the effective decay rate and the corresponding matrix elements in the diffusion matrix of the photon mode.
+Besides, the coherent feedback enlarges the parameter region supporting one-way steering and enables the transition from one-way to two-way steering.
+The frequency and amplitude ratio play an important role in enhancing and controlling the entanglement.
 Moreover, the system can sustain relatively high purity while exhibiting strong squeezing, thereby improving the quality of the local nonclassical resource state.
-Besides, we also find that compared to the system only modulated by OPA and MPA, our entanglement and steering is more robust against the thermal noises.
-Our work may provide a more controllable quantum resource, which have potential applications in continuous-variable quantum communication and quantum key distribution.
-VI Acknowledgments:
+Modulating the phase shift provides an effective way in controlling quantum correlations.
+Finally, we find that the quantum correlations in the COM system are more robust against thermal noise compared to the system only modulated by the OPA and MPA.
+Our work may provide a more controllable route to engineering quantum resource, with potential applications in continuous-variable quantum communication and quantum key distribution.
+Data availability statement
+All data that support the findings of this study are included within the article (and any supplementary files).
+Acknowledgments
 Mei Zhang thanks Junzhong Yang for helpful discussions.
@@
 Taking the expectation values of the QLEs and neglecting the zero-mean noise terms, we obtain the mean-field equation as follows,
-$\dot{\alpha}=-\bigl(\kappa_{\mathrm{fb}}+i\Delta_{\mathrm{fb}}\bigr)\alpha+ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+tE,$ (16)
+$\dot{\alpha}=-\bigl(\kappa_{\mathrm{fb}}+i\Delta_{\mathrm{fb}}\bigr)\alpha+ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+tE,$ (A1)
 where we denote $\alpha\equiv\langle a\rangle$ and $\beta\equiv\langle b\rangle$. Then we can obtain the evolution of the photon number $n_{c}(t)=|\alpha(t)|^{2}$ as
-$\frac{d}{dt}|\alpha|^{2}=\alpha^{\ast}\dot{\alpha}+\alpha\dot{\alpha}^{\ast}=-2\kappa_{\mathrm{fb}}(r)|\alpha|^{2}+2\,\mathrm{Re}\!\left[\alpha^{\ast}\Bigl(ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+t_{b}E\Bigr)\right].$ (17)
-In the long-time regime the system settles into a time-periodic limit cycle. Averaging Eq. (17) over one period $T$ of the limit cycle, we obtain the power-balance relation
-$2\kappa_{\mathrm{fb}}(r)\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}=P_{\mathrm{om}}(r)+P_{\mathrm{OPA}}(r)+P_{\mathrm{drv}},$ (18)
-where the right-hand side corresponds to the cycle-averaged power injected into the cavity by the optomechanical interaction, by the OPA and by the external drive, respectively.
+$\frac{d}{dt}|\alpha|^{2}=\alpha^{\ast}\dot{\alpha}+\alpha\dot{\alpha}^{\ast}=-2\kappa_{\mathrm{fb}}(r)|\alpha|^{2}+2\,\mathrm{Re}\!\left[\alpha^{\ast}\Bigl(ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+t_{b}E\Bigr)\right].$ (A2)
+In the long-time regime the system settles into a time-periodic limit cycle. Averaging Eq. (A2) over one period $T$ of the limit cycle, we obtain the power-balance relation
+$2\kappa_{\mathrm{fb}}(r)\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}=P_{\mathrm{om}}(r)+P_{\mathrm{OPA}}(r)+P_{\mathrm{drv}},$ (A3)
+where the right-hand side corresponds to the cycle-averaged power injected into the cavity by the optomechanical interaction, by the OPA, and by the external drive, respectively.
 For the parameter set used in Fig.
@@
 In this experimentally relevant limit Eq.
-(18) yields the approximate scaling
-$\bigl\langle|\alpha|\bigr\rangle_{T}\propto\frac{1}{\kappa_{\mathrm{fb}}(r)}=\frac{1}{\kappa(1-2r\cos\theta)},\qquad$ (19)
+(A3) yields the approximate scaling
+$\bigl\langle|\alpha|\bigr\rangle_{T}\propto\frac{1}{\kappa_{\mathrm{fb}}(r)}=\frac{1}{\kappa(1-2r\cos\theta)},$ (A4)
 which explains why both the oscillation amplitude in Fig. 2(a) and the radius of the limit-cycle trajectories in Fig. 2(b) increase monotonically with the reflectivity $r$.
@@
 To avoid confusion with the Floquet period, we denote the feedback delay time by $t_{d}$ and the loop phase by $\phi_{f}$.
-The feedback-modified input field is then written as
-$\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}(t)=r_{b}e^{\mathrm{i}\phi_{f}}a_{\mathrm{out}}(t-t_{d})+t_{b}a_{\mathrm{in}}(t).$ (20)
+The feedback modified input field is then written as
+$\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}(t)=r_{b}e^{\mathrm{i}\phi_{f}}a_{\mathrm{out}}(t-t_{d})+t_{b}a_{\mathrm{in}}(t),$ (B1)
 Using the standard input-output relation
-$a_{\mathrm{out}}(t-t_{d})=\sqrt{2\kappa_{a}}\,a(t-t_{d})-a_{\mathrm{in}}(t-t_{d}),$ (21)
+$a_{\mathrm{out}}(t-t_{d})=\sqrt{2\kappa_{a}}\,a(t-t_{d})-a_{\mathrm{in}}(t-t_{d}),$ (B2)
 the linearized Langevin equation for the cavity fluctuation becomes
-$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (22)
+$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (B3)
 $\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\,\delta a(t-t_{d})+\mathrm{noise},$
 where
-$\Delta_{a}^{\prime}=\Delta_{a}-g\bigl(\langle b\rangle+\langle b\rangle^{\ast}\bigr).$ (23)
+$\Delta_{a}^{\prime}=\Delta_{a}-g\bigl(\langle b\rangle+\langle b\rangle^{\ast}\bigr).$ (B4)
 The delay enters only through the feedback term proportional to $\delta a(t-t_{d})$.
 When the loop delay is sufficiently short, one may expand the delayed cavity fluctuation as
-$\delta a(t-t_{d})=\delta a(t)-t_{d}\,\delta\dot{a}(t)+\mathcal{O}(t_{d}^{2}).$ (24)
-Substituting Eq. (24) into Eq. (22) yields
-$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (25)
-$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\left[\delta a(t)-t_{d}\,\delta\dot{a}(t)\right]+\mathrm{noise}.$
+$\delta a(t-t_{d})=\delta a(t)-t_{d}\,\delta\dot{a}(t)+\mathcal{O}(t_{d}^{2}).$ (B5)
+Substituting Eq. (B5) into Eq. (B3) yields
+$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (B6)
+$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\left[\delta a(t)-t_{d}\,\delta\dot{a}(t)\right]+\mathrm{noise},$
 Rearranging the terms gives
-$\left[1+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}t_{d}\right]\delta\dot{a}(t)$ $=\left[-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\right]\delta a(t)$ (26)
-$\quad+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+\mathrm{noise}.$
+$\left[1+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}t_{d}\right]\delta\dot{a}(t)$ $=\left[-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\right]\delta a(t)$ (B7)
+$\quad+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+\mathrm{noise},$
 Therefore, the first-order expansion is valid provided that the prefactor of $\delta\dot{a}(t)$ remains close to unity, namely,
-$\left|2\kappa_{a}r_{b}t_{d}\right|\ll 1.$ (27)
+$\left|2\kappa_{a}r_{b}t_{d}\right|\ll 1.$ (B8)
 This condition gives the criterion under which the propagation delay can be safely neglected and the coherent feedback can be treated as instantaneous.
+ORCID iDs